# On the Encapsulation of Medical Imaging AI Algorithms

### v1

@@ 2.6 Relevant Standards
First and foremost, the DICOM standard is certainly the most important standard in this field. Embracing the DICOM standard for algorithm inputs makes it possible to associate image data with the necessary metadata for their proper interpretation, but it also serves for provenance tracking and for associating algorithm outputs with the corresponding inputs, even across multiple analysis steps and with varying types of results, such as registration matrices, segmentation masks, meshes, or measurements [fedorov_dicom_2016].
Secondly, clinical terminology systems are relevant for uniquely identifying semantic concepts that should be associated with encoded objects. For instance, instead of just assigning a segmented structure a name, ID, or label such as "kidney", one should additionally refer to at least one coding system. The DICOM standard does this in many places, for instance, allowing segmentation results to be linked to a code from the SNOMED Clinical Terms (where kidney would be identified as 64033007, together with the "SCT" coding scheme designator). While SNOMED (the use of which requires licensing) offers the largest ontology, there are a number of alternatives, often more specialized, some of which are open source, such as the Foundational Model of Anatomy (FMA [fma]) or the Terminologia Anatomica [ta2], for instance.

### v2

@@ 2.6 Relevant Standards
First and foremost, the DICOM standard is certainly the most important standard in this field. Embracing the DICOM standard for algorithm inputs makes it possible to associate image data with the necessary metadata for their proper interpretation, but it also serves for provenance tracking and for associating algorithm outputs with the corresponding inputs, even across multiple analysis steps and with varying types of results, such as registration matrices, segmentation masks, meshes, or measurements [6].
Secondly, clinical terminology systems are relevant for uniquely identifying semantic concepts that should be associated with encoded objects. For instance, instead of just assigning a segmented structure a name, ID, or label such as "kidney", one should additionally refer to at least one coding system. The DICOM standard does this in many places, for instance, allowing segmentation results to be linked to a code from the SNOMED Clinical Terms (where kidney would be identified as 64033007, together with the "SCT" coding scheme designator). While SNOMED offers the largest ontology, there are a number of alternatives, often more specialized, some of which are open source, such as the Foundational Model of Anatomy (FMA [24]) or the Terminologia Anatomica [25], for instance. Note that the use of SNOMED CT may require licensing in general, but several countries already pay nation-wide fees, supporting development and offering free use within the respective country. Furthermore, the DICOM standard contains a relevant subset of SNOMED terms that are free of use in the context of DICOM
@@
With respect to a metadata schema, we have already mentioned the FAIR4ML and FAIR4RS working groups [3, 1] which address aspects of machine learning models and research software, respectively. Image analysis algorithms are on a level in between these two, and aspects such as the inputs and outputs of algorithms, as well as their technical interfaces would require new schema concepts. To the best of our knowledge, there is no dedicated working group or standardization effort for this specific case yet.
The concept of Research Object Crates (RO-Crates) is then a relevant standardization effort that could be used to "attach" metadata to algorithm artifacts [26]. It is not specific to algorithms, but it provides a general mechanism to bundle data (in our case the executable algorithm and any AI models contained therein) and metadata in a structured way, using JSON-LD as a serialization format.